Two ingredients are used to make flat roof top filter. The first ingredient contains 35% coal tar and the second contains 60% coal tar. How much of each ingredient should be mixed to make 60 tons of roof top filter that contains 50% coal tar? Thank you
Answer by checkley77(12844)

.60x+.35(60-x)=.50*60
.60x+21-.35x=30
.25x=30-21
.25x=9
x=9/.25
x=36 tons of 60% solution is used.
60-36=24 tons of 35% solution is used.
Proof:
.60*36+.35*24=.50*60
21.6+8.4=30
30=30